OpenStudy (anonymous):

Two congruent "door-stop" trapezoids are placed edge-to-edge with measurements and markings as shown. Find the length of

OpenStudy (anonymous):

find the length of

OpenStudy (anonymous):

EB^2=BF^2-EF^2=25^2-20^2=625-400=225 EB=sqrt(225)=15 EB=AB=BC(given) so AC=AB+AB=EC+EC=15+15=30

OpenStudy (anonymous):

\[BF ^{2} = BE ^{2} + EF ^{2}\] \[( 25cm )^{2} = BE ^{2} + ( 20cm )^{2}\] BE = 15cm since BE = AB = BC, and AC = AB + BC According to the figure AC = 2*BE AC= 2*15cm AC = 30 cm